Circular No. 8734

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2006 O2 (GARRADD) G. J. Garradd reports his discovery of a comet on CCD images taken

with the 0.5-m Uppsala Schmidt telescope (discovery observation tabulated below), the object showing a broad 40′′ tail in p.a. ∼ 120^{o}.

2006 UT  α_{2000}  δ_{2000}  Mag.
July 30.38855  14^{h}44^{m}32.^{s}89  −39^{o}39′31′′.8  17.4

The available astrometry, the following parabolic orbital elements, and an ephemeris appear on MPEC 2006-P05.

T = 2006 Oct. 5.228 TT  ω = 19.^{o}660
Ω = 283.123 }2000.0
q = 1.56375 AU  i = 43.270

COMETS C/1996 R5 AND C/1996 X3–X5 (SOHO) Additional comets have been found on archival SOHO website images

(cf. IAUC 8724) — all described by K. Battams as being small and stellar in appearance. C/1996 R5, a Kreutz sungrazer, peaked at mag 7.5–8.0; the other objects belong to the Kracht group, with C/1996 X3 and C/1996 X4 peaking at mag ∼ 7.0, and C/1996 X5 peaking at mag ∼ 7.5.

Comet 1996 UT α_{2000} δ_{2000} Inst. F MPEC
C/1996 R5 Sept. 3.821 10^{h}34.^{m}6 + 6^{o}14′ C3  RK 2006-J44
C/1996 X3 Dec. 5.629 16 56.8 −22 39 C2  RK 2006-C49
C/1996 X4  5.813 16 57.0 −22 36 C2/3 RK 2006-C49
C/1996 X5  5.908 16 56.3 −22 32 C2/3 RK 2006-C49

COMETS C/1998 G9, 1998 L10, 1998 T2, 1998 W8 (SOHO) The following SOHO comets (Kreutz sungrazers except for Meyer-

group member C/1998 G9) were all small and stellar in appearance (mag 7.5–8.0), except for C/1998 L10, which was very faint and slightly diffuse.

Comet 1998 UT α_{2000} δ_{2000} Inst. F MPEC
C/1998 G9 Apr. 11.894 1^{h}22.^{m}8 +10^{o}12′ C2 RK 2006-J45
C/1998 L10 June 2.316 4 36.9 +20 17 C2 RK 2006-J45
C/1998 T2 Oct. 14.806 13 01.2 − 9 43 C3 RK 2006-J45
C/1998 W8 Nov. 30.064 16 24.5 −23 15 C3 RK 2006-J45
2006 August 1  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2006 CBAT  Daniel W. E. Green